Summary for Owl shift, July 26, 1999

Angela Bellavance (bellavan@rice.edu)
Mon, 26 Jul 1999 07:06:19 -0500 (CDT)

Shift Summary for OWL shift, 26/Jul/99

On Shift: A.Bellavance, S.Glazov (impersonated by Arun for first 2 hrs.)

Hours of beam delivered at all intensities: 6:40 / 7:00
Hours of beam delivered at nominal intensity: 6:40 / 7:00 (@3.5e12)
Hours of E832 data (special runs not included): 6:39 / 7:00
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: various aborts, energy conservations, seed
misalignment, KFREND

========================================================================
Runs ended this shift:
Run 14023 - ongoing - BEAM_832

Tapes brought to FCC this shift: QK*170-172, QKB058 (*=C,S,K)
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 datataking (were told by prev. shift: no muon runs, no magnet flip)

online kumacs:
-vvh: <adc> vs. V0 chan w/TDC hit has one channel too low = frowny
-csi: ch. 961 has comparitor probs. w/ frequency 6 at 3:00. (8 at 5am)
-muon: note to lower HV on MU3X on next access
-fpedcal: crate 10 has one channel at -35
-experts sent e-mail

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14023 00:00 start shift; taking beam; wondering where Sasha is
00:20 Arun takes tapes to Feynman.
14023 286-7 00:49 Beam aborts, returns.
14023 292-298 00:52 Noticed NM2SEED2 position was off. Called MCR and
they corrected it.
02:08 Sasha arrives after being called; didn't
realize he was on shift. (If shift list only comes
out a couple days before it goes into effect,
should people with the first few shifts be called
to make sure they know?)
14023 385-390 02:59 Combination abort and energy conservations.
14023 415 Single spill abort.
14023 426 Single spill abort.
14023 04:09 Onsplit tape mount requested (QKN015). Done.
14023 438 Single spill energy conservation period.
14023 458 Single spill energy conservation period.
14023 461 Single spill abort.
14023 528 "KTeVAna errors, please run KFREND." Done.
14023 534 06:15 Notice scaler window hasn't updated since spill
531. SDAQ program kills and restarts the window
all by itself. Wow.
14023 540 06:27 "Laser RMS too wide" audio alarm goes off.
Channel 1011 has a Stab RMS of 30.6195%. It goes
away on it's own by the next spill.

========================================================================
Follow up Notes on Previous Shift

Slow DAQ buttons:
-VolMon2 stays on constantly b/c of UCLA power supply - cont. to ignore.
-Epicure comes on occasionally complaining about NM1U, but the Epicure
window says it's fine - can ignore.

========================================================================
Items to Follow up on Next Shift

Slow DAQ buttons:
-VolMon2 stays on constantly b/c of UCLA power supply - cont. to ignore.
-Epicure comes on occasionally complaining about NM1U, but the Epicure
window says it's fine - can ignore.

Other:
-If there is a beam abort, there will sometimes be an alarm about NM2SEED2
being off. If TV is fine, ignore the alarm. However if TV is not ok,
call MCR.
-Beam will be going down at 9:30 to conserve electricity for the day.

_______
Angela Bellavance work: (630) 840-4144
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke